• Log InLog In
  • Register
Liquid`
Team Liquid Liquipedia
EDT 06:17
CEST 12:17
KST 19:17
  • Home
  • Forum
  • Calendar
  • Streams
  • Liquipedia
  • Features
  • Store
  • EPT
  • TL+
  • StarCraft 2
  • Brood War
  • Smash
  • Heroes
  • Counter-Strike
  • Overwatch
  • Liquibet
  • Fantasy StarCraft
  • TLPD
  • StarCraft 2
  • Brood War
  • Blogs
Forum Sidebar
Events/Features
News
Featured News
[ASL22] Ro4 Preview: Mirror Mirror6[ASL22] Ro8 Preview: Within Reach5[ASL22] Ro8 Preview: In A Tizzy11[ASL22] Ro16 Preview: Holy Diver5[ASL22] Ro16 Preview: Rough Waters10
Community News
Weekly Cups (October 5-11): Farewell to 5.0.162Rollback Netcode in Brood War1045.0.17 Patch Notes - October 7, 202624Weekly Cups Results (Sep 28-Oct 4)0SC4ALL: II SC2 Complete Invited Player Lineup10
StarCraft 2
General
Team Liquid Map Contest #22: Results and Winners SCII needs to put more weight on strategy and tact How do you feel about the mass reverts in the 5.0.17 PTR? Server Blocker Stellar Fest 2 Replay Pack
Tourneys
Master Swan Open (Global Bronze-Master 2) $3,500 WardiTV Not-Summer Championship Stellar Fest TWO the Moon (Dec 16-20) ISSL (IntoTheTV x SOOP SC2 League): Premier RSL 2026 Offline Finals LCQ Sign-Ups
Strategy
[H] ZvP Mid-Late Game: Stalkers Collossi HT
Custom Maps
[G] How to playtest arcade maps [A] Nemrods 1/4 players [M] (2) Sweltering Sands [M] (2) Frigid Storage
External Content
The PondCast: SC2 News & Results Mutation # 547 Temple Of Terror Mutation # 546 Catch the Train Mutation # 545 And Drops and Rifts
Brood War
General
Rollback Netcode in Brood War Bot on ladder BW General Discussion Sagi.gg Launcher Released BGH Auto Balance -> http://bghmmr.eu/
Tourneys
ASL 22 Finals 2v2v2v2 Tournament Small VOD Thread 2.0 [Women's StarCraft] MonstarZ vs Dreams 2026-10-15
Strategy
Cliff Jump Revisited (1 in a 1000 strategy) Simple Questions, Simple Answers Replay Review Process - What do you do?
Other Games
General Games
Nintendo Switch Thread Stormgate/Frost Giant Megathread Dawn of War IV General RTS Discussion Thread Warcraft III: The Frozen Throne
Dota 2
Dota 2 Champions League Season 3 Begins April 25!
League of Legends
[TL LoL EUW IHs] Teemo shall perish
Heroes of the Storm
Heroes of the Storm 2.0
Hearthstone
Deck construction bug
TL Mafia
TL Mafia Community Thread
Community
General
US Politics Mega-thread The Chess Thread Russo-Ukrainian War Thread Things Aren’t Peaceful in Palestine Artificial Intelligence Thread
Fan Clubs
Serral Fan Club
Media & Entertainment
[Manga] One Piece Movie Discussion! Diablo Animated Series on Netflix
Sports
MLB/Baseball 2023 Football (Soccer) Thread TeamLiquid Health and Fitness Initiative For 2023
World Cup 2022
Tech Support
Computer Build, Upgrade & Buying Resource Thread
TL Community
Recent Gifted Posts
Blogs
The Link Between Gaming and …
TrAiDoS
[ASL22] Ro4 Day1 Ticket Giv…
bITt.mAN
38 yo Retired SWE loo…
PurE)Rabbit-SF
Can Bots Beat Pros?? Starcr…
namkraft
[meme] I finally understa…
LUCKY_NOOB
Customize Sidebar...

Website Feedback

Closed Threads



Active: 9345 users

Brainteaser for TeamLiquid! - Page 9

Forum Index > General Forum
Post a Reply
Prev 1 7 8 9 10 11 23 Next All
qrs
Profile Blog Joined December 2007
United States3637 Posts
June 10 2011 06:53 GMT
#161
On June 10 2011 15:36 Tektos wrote:
To everyone who says 1/5 is the correct answer:

Are you assuming getting Z P is a different event as getting P Z?
Well, yes, because they are different events, aren't they?

I play 2 games as random, at least one is zerg. The possible outcomes given this constraint are:

I play 2 games as zerg
I play 1 game as zerg and 1 as protoss
I play 1 game as zerg and 1 as terran


3 possible outcomes, playing zerg twice is one of the possible outcomes:
1 / 3

I play 2 games as Zerg
I play game 1 as Zerg and game 2 as Protoss.
I play game 1 as Protoss and game 2 as Zerg.
I play game 1 as Zerg and game 2 as Terran.
I play game 1 as Terran and game 2 as Zerg.

5 possible outcomes; playing Zerg twice is one of the possible outcomes.

Why aren't you counting ZP and PZ separately?
'As per the American Heart Association, the beat of the Bee Gees song "Stayin' Alive" provides an ideal rhythm in terms of beats per minute to use for hands-only CPR. One can also hum Queen's "Another One Bites The Dust".' —Wikipedia
BlackJack
Profile Blog Joined June 2003
United States10574 Posts
June 10 2011 06:54 GMT
#162
Here's a similar problem that my stats teacher gave me some years ago. It's something along these lines.

You are at a party chatting with a guest and they share that they have 2 children. Right after they tell you this a girl walks up to the person you are talking to and the person says "oh, here is my daughter now." The question is what is the probability that the other child is also a girl. The answer I was given was 1/3 for the same reason, 3 possibilities, BG, GB, and GG so 1 out of 3, right?

Here's my problem.. The fact that the girl walked up is a completely random occurrence. It could have just as easily been a boy. If it was a boy, you would also say that the chance of the person having 2 boys is 1/3. Now you must be saying that the probability the unknown offspring is the opposite gender of the first offspring you see is 2/3. We know that the probability of having 2 different gender children vs 2 same-sex children is 50-50. What am I missing?
darmousseh
Profile Blog Joined May 2010
United States3437 Posts
June 10 2011 06:55 GMT
#163
This is very simple as many others have said. Given that I played either game 1 as zerg or game 2 as zerg, what are the odds that i played zerg two times in a row.

First off, there are 9 possibile combinations pp pz pt tp tz tt zp zz zt and only in 1 scenario is zz. Get rid of any scenarios in which you don't play zerg and there are 5 left, thus 1/5.

There are many ways to do this problem and they all result in 1/5.
Developer for http://mtgfiddle.com
qrs
Profile Blog Joined December 2007
United States3637 Posts
Last Edited: 2011-06-10 06:57:53
June 10 2011 06:57 GMT
#164
On June 10 2011 15:51 Tektos wrote:
Show nested quote +
On June 10 2011 15:47 OneOther wrote:
Haha wow there seems to be a lot of debate over a simple conditional probability problem. Yes, the wording was a little bit unclear, but his edits clarified what he was talking about. Some people seem to just not understand the concept of a conditional probability, or are still misreading the problem.

"I played two games as random today. In at least one of the games I was Zerg. (NOT the first game, but just in one of the two - this is an important distinction that the OP was a little bit unclear on at first but wasn't necessarily wrong either). Given this, what is the probability that I played both games as Zerg? (Not what the probability of getting Zerg in the second game was)"

Conditional probability P (A l B) = P(A and B) / P(A)
(Probability of getting Zerg in two games given Zerg in at least one game) / (Probability of getting Zerg in one game)

The possible combinations are ZZ, ZP, ZT, TZ, PZ, PP, TT, TP, PT.
Probability of getting Zerg in one game = 5/9 (since PP, TT, TP and PT are eliminated)
Given this probability, the probability of getting another Zerg = 1/9 (only ZZ remains out of the 9)

Therefore, using the equation, we arrive at 1/9 divided by 5/9 = 9/40, or 1/5.


ZP and PZ are the same outcome (playing once as zerg and once as protoss)
TZ and ZT are the same outcome (playing once as zerg and once as terran)

That consolidates your possible outcomes down to 3 results, giving 1/3 chance that the other game will be zerg.
By your reasoning, I can prove to you that you have a fifty percent chance of rolling a 1, using a fair die.

There are two possibilities: you roll 1 or you don't roll 1. 2, 3, 4, 5, 6 are the same result: not 1. That consolidates your possible outcomes down to 2 results, giving 1/2 chance that you roll a 1.

You see the flaw in this, don't you?
'As per the American Heart Association, the beat of the Bee Gees song "Stayin' Alive" provides an ideal rhythm in terms of beats per minute to use for hands-only CPR. One can also hum Queen's "Another One Bites The Dust".' —Wikipedia
Tektos
Profile Joined November 2010
Australia1321 Posts
June 10 2011 06:57 GMT
#165
On June 10 2011 15:52 Stropheum wrote:
Since the probability of getting zerg in 1 game is 1/3, getting zerg in 2 games = 2/6, which simplifies to 1/3 obviously, but is needed for the next step.
Since you played 1 game already and got zerg 1 time, your probability for 2 zerg games remains 2/6, yet you factor in that you've already played one game, reducing it to 2/5, and you factor in the fact that you've gotten zerg, making the probability 1/5 that you will get zerg a second time.
Do i win?


No the probability of getting zerg in two consecutive games is 1/3 * 1/3 = 1/9
EmeraldSparks
Profile Blog Joined January 2008
United States1451 Posts
June 10 2011 06:58 GMT
#166
On June 10 2011 15:17 Tektos wrote:
Show nested quote +
On June 10 2011 15:12 EmeraldSparks wrote:
On June 10 2011 10:11 theDreamStick wrote:
It is correct that I am asking a conditional probability question:
Given that at least one of my games was Zerg, what is the probability that both of my games are zerg.

It's either 100%, or 0%, depending on whether or not both games were zerg.

I'll give you a puzzle: I either ate a sandwich today, or I didn't.

What is the probability that I ate a sandwich?


With the information provided nobody is able to tell you what the probability that you eat a sandwich is, as eating a sandwich vs. not eating a sandwich is not a fixed probability event. Some people eat sandwiches every day, some people eat sandwiches once a week, some people never eat sandwiches. The only information that can be presented is that the probability that it is either of those two events is 100% because the probability of an event added to the compliment of that event is always 100%.

You could, however, give us a various sample of consecutive days stating whether you ate a sandwich on that day or not and an approximate probability to a certain confidence level based on sample size could be predicted provided the event of eating a sandwich is not pattern based.

So it's based on a huge set of assumptions and extra information.

Let's say that it happened like this:

Your friend sits down to watch TLO play a game. He randoms.

He only manages to watch one of the games.

After the game, he tells you, "holy shit, TLO's zerg is fucking awesome."

Or "TLO's protoss is fucking awesome."

Or "TLO's terran is fucking awesome."

Then, he asks you, because he's a clever sadistic bastard,

"What's the chance that TLO was zerg twice?"

Or "What's the chance that TLO was protoss twice?"

Or "What's the chance that TLO was terran twice?"

The answer to the first question is one third.

The answer to the second question is one third.

The answer to the third question is one third.

So the answer is one third if the backstory happened like this. But it's 1/5th if the backstory happened another way, say that your friend is this disgustingly patriotic zerg fanboy.and would only tell you that TLO played a zerg game and he doesn't give a shit about terran or protoss games.

But since you don't know the backstory, you can't answer the question.

The same way you can't answer my question without knowing if I ate a sandwich today.

Solution: + Show Spoiler +
I haven't had a sandwich in weeks.
But why?
OneOther
Profile Blog Joined August 2004
United States10774 Posts
June 10 2011 06:58 GMT
#167
On June 10 2011 15:51 Tektos wrote:
Show nested quote +
On June 10 2011 15:47 OneOther wrote:
Haha wow there seems to be a lot of debate over a simple conditional probability problem. Yes, the wording was a little bit unclear, but his edits clarified what he was talking about. Some people seem to just not understand the concept of a conditional probability, or are still misreading the problem.

"I played two games as random today. In at least one of the games I was Zerg. (NOT the first game, but just in one of the two - this is an important distinction that the OP was a little bit unclear on at first but wasn't necessarily wrong either). Given this, what is the probability that I played both games as Zerg? (Not what the probability of getting Zerg in the second game was)"

Conditional probability P (A l B) = P(A and B) / P(A)
(Probability of getting Zerg in two games given Zerg in at least one game) / (Probability of getting Zerg in one game)

The possible combinations are ZZ, ZP, ZT, TZ, PZ, PP, TT, TP, PT.
Probability of getting Zerg in one game = 5/9 (since PP, TT, TP and PT are eliminated)
Given this probability, the probability of getting another Zerg = 1/9 (only ZZ remains out of the 9)

Therefore, using the equation, we arrive at 1/9 divided by 5/9 = 9/40, or 1/5.


ZP and PZ are the same outcome (playing once as zerg and once as protoss)
TZ and ZT are the same outcome (playing once as zerg and once as terran)

That consolidates your possible outcomes down to 3 results, giving 1/3 chance that the other game will be zerg.

ZP and PZ are different outcomes. TZ and ZT are different outcomes.
Getting a Zerg in the first game and Protoss in the second game is different from getting Protoss in the first game and then Zerg in the second. Same for ZT/TZ.


On June 10 2011 15:51 ZombiesOMG wrote:
Show nested quote +
On June 10 2011 15:47 OneOther wrote:

The possible combinations are ZZ, ZP, ZT, TZ, PZ, PP, TT, TP, PT.
Probability of getting Zerg in one game = 5/9 (since PP, TT, TP and PT are eliminated)
Given this condition, the probability of getting another Zerg = 1/9 (only ZZ remains out of the 9)

Therefore, using the equation, we arrive at 1/9 divided by 5/9 = 9/45, or 1/5.


I just don't understand how ZP and PZ, or TZ and ZT are different. The OP can only be one of 3 races, why does the opponent's race have any bearing at all? Doesn't that invalidate this calculation?

What do you mean? The opponent's races don't have to do anything with the problem.
Tektos
Profile Joined November 2010
Australia1321 Posts
June 10 2011 06:59 GMT
#168
On June 10 2011 15:57 qrs wrote:
Show nested quote +
On June 10 2011 15:51 Tektos wrote:
On June 10 2011 15:47 OneOther wrote:
Haha wow there seems to be a lot of debate over a simple conditional probability problem. Yes, the wording was a little bit unclear, but his edits clarified what he was talking about. Some people seem to just not understand the concept of a conditional probability, or are still misreading the problem.

"I played two games as random today. In at least one of the games I was Zerg. (NOT the first game, but just in one of the two - this is an important distinction that the OP was a little bit unclear on at first but wasn't necessarily wrong either). Given this, what is the probability that I played both games as Zerg? (Not what the probability of getting Zerg in the second game was)"

Conditional probability P (A l B) = P(A and B) / P(A)
(Probability of getting Zerg in two games given Zerg in at least one game) / (Probability of getting Zerg in one game)

The possible combinations are ZZ, ZP, ZT, TZ, PZ, PP, TT, TP, PT.
Probability of getting Zerg in one game = 5/9 (since PP, TT, TP and PT are eliminated)
Given this probability, the probability of getting another Zerg = 1/9 (only ZZ remains out of the 9)

Therefore, using the equation, we arrive at 1/9 divided by 5/9 = 9/40, or 1/5.


ZP and PZ are the same outcome (playing once as zerg and once as protoss)
TZ and ZT are the same outcome (playing once as zerg and once as terran)

That consolidates your possible outcomes down to 3 results, giving 1/3 chance that the other game will be zerg.
By your reasoning, I can prove to you that you have a fifty percent chance of rolling a 1, using a fair die.

There are two possibilities: you roll 1 or you don't roll 1. 2, 3, 4, 5, 6 are the same result: not 1. That consolidates your possible outcomes down to 2 results, giving 1/2 chance that you roll a 1.

You see the flaw in this, don't you?


ZP vs. PZ is comepletely different from P(A) vs. P(not A) you seem to have no clue what you're doing.
darmousseh
Profile Blog Joined May 2010
United States3437 Posts
June 10 2011 06:59 GMT
#169
The correct way to phrase the problem is

"Given that I played 2 games and that at least one of them was played as zerg, what is the probability that both were played as zerg?"
Developer for http://mtgfiddle.com
EmeraldSparks
Profile Blog Joined January 2008
United States1451 Posts
June 10 2011 07:00 GMT
#170
On June 10 2011 15:59 darmousseh wrote:
The correct way to phrase the problem is

"Given that I played 2 games and that at least one of them was played as zerg, what is the probability that both were played as zerg?"

100%, if you're Idra.
But why?
Tektos
Profile Joined November 2010
Australia1321 Posts
June 10 2011 07:01 GMT
#171
On June 10 2011 15:58 OneOther wrote:
Show nested quote +
On June 10 2011 15:51 Tektos wrote:
On June 10 2011 15:47 OneOther wrote:
Haha wow there seems to be a lot of debate over a simple conditional probability problem. Yes, the wording was a little bit unclear, but his edits clarified what he was talking about. Some people seem to just not understand the concept of a conditional probability, or are still misreading the problem.

"I played two games as random today. In at least one of the games I was Zerg. (NOT the first game, but just in one of the two - this is an important distinction that the OP was a little bit unclear on at first but wasn't necessarily wrong either). Given this, what is the probability that I played both games as Zerg? (Not what the probability of getting Zerg in the second game was)"

Conditional probability P (A l B) = P(A and B) / P(A)
(Probability of getting Zerg in two games given Zerg in at least one game) / (Probability of getting Zerg in one game)

The possible combinations are ZZ, ZP, ZT, TZ, PZ, PP, TT, TP, PT.
Probability of getting Zerg in one game = 5/9 (since PP, TT, TP and PT are eliminated)
Given this probability, the probability of getting another Zerg = 1/9 (only ZZ remains out of the 9)

Therefore, using the equation, we arrive at 1/9 divided by 5/9 = 9/40, or 1/5.


ZP and PZ are the same outcome (playing once as zerg and once as protoss)
TZ and ZT are the same outcome (playing once as zerg and once as terran)

That consolidates your possible outcomes down to 3 results, giving 1/3 chance that the other game will be zerg.

ZP and PZ are different outcomes. TZ and ZT are different outcomes.
Getting a Zerg in the first game and Protoss in the second game is different from getting Protoss in the first game and then Zerg in the second. Same for ZT/TZ.



No they are the same event because the games are independent. Everyone is confusing dependent event conditional probability with independent event conditional probability.
darmousseh
Profile Blog Joined May 2010
United States3437 Posts
June 10 2011 07:01 GMT
#172
On June 10 2011 16:00 EmeraldSparks wrote:
Show nested quote +
On June 10 2011 15:59 darmousseh wrote:
The correct way to phrase the problem is

"Given that I played 2 games and that at least one of them was played as zerg, what is the probability that both were played as zerg?"

100%, if you're Idra.

I've seen idra offrace as protoss......
Developer for http://mtgfiddle.com
EmeraldSparks
Profile Blog Joined January 2008
United States1451 Posts
June 10 2011 07:02 GMT
#173
On June 10 2011 16:01 darmousseh wrote:
Show nested quote +
On June 10 2011 16:00 EmeraldSparks wrote:
On June 10 2011 15:59 darmousseh wrote:
The correct way to phrase the problem is

"Given that I played 2 games and that at least one of them was played as zerg, what is the probability that both were played as zerg?"

100%, if you're Idra.

I've seen idra offrace as protoss......

Looks like somebody made a bad assumption.

That somebody's me.
But why?
qrs
Profile Blog Joined December 2007
United States3637 Posts
June 10 2011 07:03 GMT
#174
On June 10 2011 15:59 Tektos wrote:
Show nested quote +
On June 10 2011 15:57 qrs wrote:
On June 10 2011 15:51 Tektos wrote:
On June 10 2011 15:47 OneOther wrote:
Haha wow there seems to be a lot of debate over a simple conditional probability problem. Yes, the wording was a little bit unclear, but his edits clarified what he was talking about. Some people seem to just not understand the concept of a conditional probability, or are still misreading the problem.

"I played two games as random today. In at least one of the games I was Zerg. (NOT the first game, but just in one of the two - this is an important distinction that the OP was a little bit unclear on at first but wasn't necessarily wrong either). Given this, what is the probability that I played both games as Zerg? (Not what the probability of getting Zerg in the second game was)"

Conditional probability P (A l B) = P(A and B) / P(A)
(Probability of getting Zerg in two games given Zerg in at least one game) / (Probability of getting Zerg in one game)

The possible combinations are ZZ, ZP, ZT, TZ, PZ, PP, TT, TP, PT.
Probability of getting Zerg in one game = 5/9 (since PP, TT, TP and PT are eliminated)
Given this probability, the probability of getting another Zerg = 1/9 (only ZZ remains out of the 9)

Therefore, using the equation, we arrive at 1/9 divided by 5/9 = 9/40, or 1/5.


ZP and PZ are the same outcome (playing once as zerg and once as protoss)
TZ and ZT are the same outcome (playing once as zerg and once as terran)

That consolidates your possible outcomes down to 3 results, giving 1/3 chance that the other game will be zerg.
By your reasoning, I can prove to you that you have a fifty percent chance of rolling a 1, using a fair die.

There are two possibilities: you roll 1 or you don't roll 1. 2, 3, 4, 5, 6 are the same result: not 1. That consolidates your possible outcomes down to 2 results, giving 1/2 chance that you roll a 1.

You see the flaw in this, don't you?


ZP vs. PZ is comepletely different from P(A) vs. P(not A) you seem to have no clue what you're doing.
What I'm doing is showing how your "consolidation process" can be abused 6 ways till Sunday. You're the one who doesn't seem to know what you're doing.

Playing a Zerg in game 1 and playing a Zerg in game 2 are different, mutually exclusive events. Saying that every event that includes a Zerg is the same event, is precisely as (in)valid as saying that every event that is non-1 is the same event.
'As per the American Heart Association, the beat of the Bee Gees song "Stayin' Alive" provides an ideal rhythm in terms of beats per minute to use for hands-only CPR. One can also hum Queen's "Another One Bites The Dust".' —Wikipedia
Tektos
Profile Joined November 2010
Australia1321 Posts
June 10 2011 07:06 GMT
#175
On June 10 2011 16:03 qrs wrote:
Show nested quote +
On June 10 2011 15:59 Tektos wrote:
On June 10 2011 15:57 qrs wrote:
On June 10 2011 15:51 Tektos wrote:
On June 10 2011 15:47 OneOther wrote:
Haha wow there seems to be a lot of debate over a simple conditional probability problem. Yes, the wording was a little bit unclear, but his edits clarified what he was talking about. Some people seem to just not understand the concept of a conditional probability, or are still misreading the problem.

"I played two games as random today. In at least one of the games I was Zerg. (NOT the first game, but just in one of the two - this is an important distinction that the OP was a little bit unclear on at first but wasn't necessarily wrong either). Given this, what is the probability that I played both games as Zerg? (Not what the probability of getting Zerg in the second game was)"

Conditional probability P (A l B) = P(A and B) / P(A)
(Probability of getting Zerg in two games given Zerg in at least one game) / (Probability of getting Zerg in one game)

The possible combinations are ZZ, ZP, ZT, TZ, PZ, PP, TT, TP, PT.
Probability of getting Zerg in one game = 5/9 (since PP, TT, TP and PT are eliminated)
Given this probability, the probability of getting another Zerg = 1/9 (only ZZ remains out of the 9)

Therefore, using the equation, we arrive at 1/9 divided by 5/9 = 9/40, or 1/5.


ZP and PZ are the same outcome (playing once as zerg and once as protoss)
TZ and ZT are the same outcome (playing once as zerg and once as terran)

That consolidates your possible outcomes down to 3 results, giving 1/3 chance that the other game will be zerg.
By your reasoning, I can prove to you that you have a fifty percent chance of rolling a 1, using a fair die.

There are two possibilities: you roll 1 or you don't roll 1. 2, 3, 4, 5, 6 are the same result: not 1. That consolidates your possible outcomes down to 2 results, giving 1/2 chance that you roll a 1.

You see the flaw in this, don't you?


ZP vs. PZ is comepletely different from P(A) vs. P(not A) you seem to have no clue what you're doing.
What I'm doing is showing how your "consolidation process" can be abused 6 ways till Sunday. You're the one who doesn't seem to know what you're doing.

Playing a Zerg in game 1 and playing a Zerg in game 2 are different, mutually exclusive events. Saying that every event that includes a Zerg is the same event, is precisely as (in)valid as saying that every event that is non-1 is the same event.


No it can't be abused 6 ways till Sunday unless you have no clue about a thing to do with probability. Order of occurrence does not play a factor in independent event conditional probability.
EmeraldSparks
Profile Blog Joined January 2008
United States1451 Posts
June 10 2011 07:07 GMT
#176
ZZ - Your friend reports, "TLO played at least one game as zerg."
ZP - Your friend reports, "TLO played at least one game as zerg."
ZT - Your friend reports, "TLO played at least one game as zerg."
PZ - Your friend reports, "TLO played at least one game as zerg."
TZ - Your friend reports, "TLO played at least one game as zerg."
TT - Your friend goes to the pub.
TP - Your friend goes to the pub.
PT - Your friend goes to the pub.
PP - Your friend goes to the pub.

Your friend asks you, "what is the chance that TLO played both games with the same race."

The answer to the question in this case is 1/5.

ZZ - Your friend reports, "TLO played at least one game as zerg."
ZP - Your friend reports with 50% chance "TLO played at least one game as zerg" and 50% chance "TLO played at least one game as protoss."
ZT - Your friend reports with 50% chance "TLO played at least one game as zerg" and 50% chance "TLO played at least one game as terran."
PZ - Your friend reports with 50% chance "TLO played at least one game as zerg" and 50% chance "TLO played at least one game as protoss."
TZ - Your friend reports with 50% chance "TLO played at least one game as zerg" and 50% chance "TLO played at least one game as terran."
TT - Your friend reports, "TLO played at least one game as terran."
TP - Your friend reports with 50% chance "TLO played at least one game as protoss" and 50% chance "TLO played at least one game as terran."
PT - Your friend reports with 50% chance "TLO played at least one game as protoss" and 50% chance "TLO played at least one game as terran."
PP - Your friend reports, "TLO played at least one game as protoss."

Your friend asks you, "what is the chance that TLO played both games with the same race."

The answer to the question in this case is 1/3.

You have no idea of knowing your friend's method of coming up with the question.

You cannot answer the question without further information.
But why?
qrs
Profile Blog Joined December 2007
United States3637 Posts
June 10 2011 07:08 GMT
#177
On June 10 2011 16:01 Tektos wrote:
No they are the same event because the games are independent. Everyone is confusing dependent event conditional probability with independent event conditional probability.
OK, let's do it this way. Imagine that a bunch of random games are played.

1/9 of them will be ZZ.
2/9 of them will be ZP/PZ
2/9 of them will be ZT/TZ

2/9 of them will be TP/PT
1/9 of them will be PP
1/9 of them will be TT.

Do you agree with the above?

If so, then consider the set of sets of games where at least one Z is played: 1/5 of these are ZZ.
'As per the American Heart Association, the beat of the Bee Gees song "Stayin' Alive" provides an ideal rhythm in terms of beats per minute to use for hands-only CPR. One can also hum Queen's "Another One Bites The Dust".' —Wikipedia
DeltruS
Profile Blog Joined April 2010
Canada2214 Posts
June 10 2011 07:09 GMT
#178
Thank you OneOther. Starcraft themed math makes the mundane fun.
http://grooveshark.com/#/deltrus/music
Ivs
Profile Joined January 2008
Australia139 Posts
June 10 2011 07:10 GMT
#179
On June 10 2011 16:06 Tektos wrote:
Show nested quote +
On June 10 2011 16:03 qrs wrote:
On June 10 2011 15:59 Tektos wrote:
On June 10 2011 15:57 qrs wrote:
On June 10 2011 15:51 Tektos wrote:
On June 10 2011 15:47 OneOther wrote:
Haha wow there seems to be a lot of debate over a simple conditional probability problem. Yes, the wording was a little bit unclear, but his edits clarified what he was talking about. Some people seem to just not understand the concept of a conditional probability, or are still misreading the problem.

"I played two games as random today. In at least one of the games I was Zerg. (NOT the first game, but just in one of the two - this is an important distinction that the OP was a little bit unclear on at first but wasn't necessarily wrong either). Given this, what is the probability that I played both games as Zerg? (Not what the probability of getting Zerg in the second game was)"

Conditional probability P (A l B) = P(A and B) / P(A)
(Probability of getting Zerg in two games given Zerg in at least one game) / (Probability of getting Zerg in one game)

The possible combinations are ZZ, ZP, ZT, TZ, PZ, PP, TT, TP, PT.
Probability of getting Zerg in one game = 5/9 (since PP, TT, TP and PT are eliminated)
Given this probability, the probability of getting another Zerg = 1/9 (only ZZ remains out of the 9)

Therefore, using the equation, we arrive at 1/9 divided by 5/9 = 9/40, or 1/5.


ZP and PZ are the same outcome (playing once as zerg and once as protoss)
TZ and ZT are the same outcome (playing once as zerg and once as terran)

That consolidates your possible outcomes down to 3 results, giving 1/3 chance that the other game will be zerg.
By your reasoning, I can prove to you that you have a fifty percent chance of rolling a 1, using a fair die.

There are two possibilities: you roll 1 or you don't roll 1. 2, 3, 4, 5, 6 are the same result: not 1. That consolidates your possible outcomes down to 2 results, giving 1/2 chance that you roll a 1.

You see the flaw in this, don't you?


ZP vs. PZ is comepletely different from P(A) vs. P(not A) you seem to have no clue what you're doing.
What I'm doing is showing how your "consolidation process" can be abused 6 ways till Sunday. You're the one who doesn't seem to know what you're doing.

Playing a Zerg in game 1 and playing a Zerg in game 2 are different, mutually exclusive events. Saying that every event that includes a Zerg is the same event, is precisely as (in)valid as saying that every event that is non-1 is the same event.


No it can't be abused 6 ways till Sunday unless you have no clue about a thing to do with probability. Order of occurrence does not play a factor in independent event conditional probability.


By your logic, the chance of flipping two coins and get two heads in a row is a 1/3, because the possible outcomes are HH, TT, and HT (which is the same as TH).
Abenson
Profile Blog Joined December 2009
Canada4122 Posts
June 10 2011 07:11 GMT
#180
erp.
1/5 - i don't see how hard this can be lol
All you have to do, if you're confused, is to list out all the possibilities.

As I'm posting right now, 61% have answered 1/3 lol
Prev 1 7 8 9 10 11 23 Next All
Please log in or register to reply.
Live Events Refresh
Next event in 43m
[ Submit Event ]
Live Streams
Refresh
StarCraft 2
OGKoka 1
StarCraft: Brood War
Calm 4126
Bisu 925
Horang2 564
Jaedong 269
Larva 264
BeSt 232
ZerO 215
Killer 170
Stork 154
Dewaltoss 127
[ Show more ]
Leta 124
actioN 107
Free 98
Hyuk 93
Barracks 76
Sharp 71
Mind 69
Shinee 64
ggaemo 53
Hyun 49
Pusan 47
FanTaSy 46
Hm[arnc] 40
Light 34
EffOrt 32
ToSsGirL 27
soO 24
Bale 22
sorry 19
ajuk12(nOOB) 16
Yoon 14
Noble 11
sSak 9
Counter-Strike
olofmeister1271
shoxiejesuss1075
kennyS383
Other Games
gofns31050
singsing1326
Coldzera 745
Pyrionflax134
Lowko90
ZerO(Twitch)8
BEARDiaguz2
Organizations
StarCraft: Brood War
UltimateBattle 27
[ Show 11 non-featured ]
StarCraft 2
• LUISG 32
• AfreecaTV YouTube
• intothetv
• Kozan
• IndyKCrew
• Migwel
StarCraft: Brood War
• BSLYoutube
• STPLYoutube
• ZZZeroYoutube
Dota 2
• lizZardDota2223
League of Legends
• Stunt437
Upcoming Events
Online Event
43m
WardiTV Invitational
43m
OGKoka 1
Korean StarCraft League
16h 43m
Afreeca Starleague
19h 43m
Rush vs Soulkey
CranKy Ducklings
23h 43m
WardiTV Invitational
1d
AI Arena Tournament
1d 5h
BSL Open Qualifier
1d 8h
Sparkling Tuna Cup
1d 23h
OSC
2 days
[ Show More ]
IPSL
2 days
BSL Open Qualifier
2 days
WardiTV Weekly
3 days
IntoTheTV X SOOP
4 days
PiGosaur Cup
4 days
The PondCast
4 days
Kung Fu Cup
5 days
IntoTheTV X SOOP
6 days
Liquipedia Results

Completed

Super Anchor Qualifying S3
Blizzard Classic Cup 2026
Copium Cup

Ongoing

ASL Season 22
Acropolis #5
HCC Season 3
Thunderpick World Champ. '26
ESL Pro League Season 24
Stake Ranked Episode 4
1win Private Club #1
Logitech G Play Connect 2026
SL StarSeries Fall 2026
FISSURE Playground #3
BLAST Open Fall 2026
Esports World Cup 2026

Upcoming

Acropolis #5 - GSC
BSL Season 23
SC4ALL II: Brood War
BSL 23: Non-Korean Championship
HSC XXX
Stellar Fest 2: Lunar Cup
SC4ALL II: StarCraft II
Kung Fu Cup 2026 Grand Finals
RSL Offline Finals
eXTREMESLAND 2026
PGL Major Singapore 2026
Stake Ranked Episode 6
BLAST Rivals Fall 2026
IEM Beijing 2026
Stake Ranked Episode 5
PGL Masters Bucharest 2026
1win Private Club #2
TLPD

1. ByuN
2. TY
3. Dark
4. Solar
5. Stats
6. Nerchio
7. sOs
8. soO
9. INnoVation
10. Elazer
1. Rain
2. Flash
3. EffOrt
4. Last
5. Bisu
6. Soulkey
7. Mini
8. Sharp
Sidebar Settings...

Advertising | Privacy Policy | Terms Of Use | Contact Us

Original banner artwork: Jim Warren
The contents of this webpage are copyright © 2026 TLnet. All Rights Reserved.